Question:
If the probability of winning a game is 0.3, the probability of losing it is
यदि किसी खेल को जीतने की प्रायिकता 0.3 है, तो इसे हारने की प्रायिकता क्या है?
Options:
- 0.9
- 0.2
- 0.7
- 0.5
- 0.9
- 0.2
- 0.7
- 0.5
Correct Answer: c)
Explanation:
\(P(\text{winning}) + P(\text{losing}) = 1 \implies P(\text{losing}) = 1 – 0.3 = 0.7\).
\(P(\text{जीतना}) + P(\text{हारना}) = 1 \implies P(\text{हारना}) = 1 – 0.3 = 0.7\)।